OpenStudy (anonymous):

at time t=0, an object is observed at x=0; and its position along the x-axis follows this expression: x=-3t+t cubed, where the units for distance and time are meters and seconds, respectively. What is the object's displacement delta x between t=1.0s and t=3.0s?

OpenStudy (anonymous):

could you help me out with a problem?

OpenStudy (anonymous):

I am having trouble understanding this one

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

x=-3t+t x=-3*0+0

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

where did you get 0?

OpenStudy (anonymous):

from 0 seconds?

OpenStudy (anonymous):

X=0 and T=0 so in the problem X=-3T+T so you just take 0 and plug it in to T and 0and plug it in to X

OpenStudy (anonymous):

ok, I got 0 for x... now what?

OpenStudy (anonymous):

the final answer is +20m

OpenStudy (anonymous):

I am trying to figure out the object's displacement between t=1.0s and t=3.0s

OpenStudy (anonymous):

i just got told by my teacher to plug the numbers that are at the boddom of the problem

OpenStudy (anonymous):

hmmm..

OpenStudy (anonymous):

4=-3*1+1 36=-3*3+3 4 and 36

OpenStudy (anonymous):

9+27=36